Calculate IV rate for the pump:

1. The order: 900 ml LR IV 7am – 7pm. How many ml/hr do we set the pump to?

2.The order: 1000 ml NS to infuse over 3 hours. How many ml/hr would we set the pump

to?

3. The order: Infuse 750 ml D5W from 6:00am – 11:00am. How many ml/hr would this

be?
Answer #1

Ans) Formula:

ml/hr = Total volume to be administered ÷ time in hrs

1) The order: 900 ml LR IV 7am – 7pm. i.e 12 hours

= 900÷ 12 = 75 ml/hr.

2) The order: 1000 ml NS to infuse over 3 hours

= 1000÷ 3 = 333.33 ~ 333 ml/hr.

3) The order: Infuse 750 ml D5W from 6:00am – 11:00am i.e 5 hrs

= 750 ÷ 5 = 150 ml/ hr